40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL, Page 52 of 52  Not logged in ELOG logo
ID Date Author Type Category Subject
  42   Tue Feb 2 16:36:46 2010 FrankElectronicsDAQDAQ pinouts

3123-card (16bit input), 25pin d-sub connector

channel LO HI
0 2 14
1 16 3
2 5 17
3 19 6
4 8 20
5 22 9
6 11 23
7 25 12

 


 4116-card (16bit output), 50-pin connector

channel LO HI
0 3 4
1 5 6
2 7 8
3 9 10
4 11 12
5 13 14
6 15 16
7 17 18

 

  41   Mon Feb 1 12:51:43 2010 Rana, FrankDailyProgressNoiseBudgetFirst Refcav noise measurement
Attached is the relative frequency noise between the 2 reference cavities downstairs.

The blue curve shows the calibrated control signal from the PDH loop of the analyzer cavity adjusting the AOM frequency.

The low frequency hump is caused by pointing introduced by the AOM used to lock the second (analyzer) cavity to the first (reference cavity).
Changing the frequency causes causes beam-pointing which, at the moment, is much larger as expected. This might be due to bad alignment of the double-passing beam.

The flat part seems to be the shot noise level of the light to read the analyzer cavity.
We have to increase the power in this part of the setup but have to replace a mirror for that and re-align everything (all EOMs, AOM, cavities etc. That's why we waited until the first result.

Current total power in that part is 1.8mW. The other cavity uses 14.7mW.

The bump at a couple of kHz is due to the PDH loop of the analyzer cavity; phase margin is not enough, but we use one of the unmodified generic PDH boxes so far so it's not a surprise.

Green curve shows the calibrated error point signal. The overall noise performance is ~1000x worse than it should be.
Attachment 1: refcav-noise.pdf
refcav-noise.pdf
  40   Thu Jan 28 16:07:29 2010 FrankComputingDAQpsl channels moved to C3 + new channels

here is a list of all channels of the psl subsystem. We changed the generic channel names to final names now.
Refcav channels are now C3:PSL-RCAV and analyzer cavity channels are C3:PSL-ACAV. Rest see below...

#####################
# 10W MOPA channels #
#####################
[C3:PSL-126MOPA_AMPMON]        # internal laser power monitor
[C3:PSL-126MOPA_126MON]        # internal NPRO power monitor
[C3:PSL-126MOPA_DS1]            # diode sensor 1
[C3:PSL-126MOPA_DS2]            # diode sensor 2
[C3:PSL-126MOPA_DS3]            # diode sensor 3
[C3:PSL-126MOPA_DS4]            # diode sensor 4
[C3:PSL-126MOPA_DS5]            # diode sensor 5
[C3:PSL-126MOPA_DS6]            # diode sensor 6
[C3:PSL-126MOPA_DS7]            # diode sensor 7
[C3:PSL-126MOPA_DS8]            # diode sensor 8
[C3:PSL-126MOPA_126PWR]        # NPRO power monitor
[C3:PSL-126MOPA_DTMP]        # diode temperature
[C3:PSL-126MOPA_LTMP]        # pump diode temperature
[C3:PSL-126MOPA_DMON]        # diode output monitor
[C3:PSL-126MOPA_LMON]        # pump diode output monitor
[C3:PSL-126MOPA_CURMON]        # pump diode current monitor
[C3:PSL-126MOPA_DTEC]        # diode heater voltage
[C3:PSL-126MOPA_LTEC]        # pump diode heater voltage
[C3:PSL-126MOPA_CURMON2]        # pump diode current monitor
[C3:PSL-126MOPA_HTEMP]        # head temperature
[C3:PSL-126MOPA_HTEMPSET]        # head temperature set point
[C3:PSL-126MOPA_FAULT]        # laser fault indicator
[C3:PSL-126MOPA_INTERLOCK]        # interlock control
[C3:PSL-126MOPA_SHUTTER]        # shutter control
[C3:PSL-126MOPA_126LASE]        # NPRO lase status
[C3:PSL-126MOPA_AMPON]        # power amplifier lase status
[C3:PSL-126MOPA_SHUTOPENEX]        #
[C3:PSL-126MOPA_STANDBY]        #
[C3:PSL-126MOPA_126NE]        # NPRO noise eater
[C3:PSL-126MOPA_126STANDBY]        # NPRO standby
[C3:PSL-126MOPA_DCAMP]        #
[C3:PSL-126MOPA_126CURADJ]        #
[C3:PSL-126MOPA_126SLOW]        #
[C3:PSL-126MOPA_BEAMON]        # beam on logical

#######################
# 80 MHz VCO channels #
#######################
[C3:PSL-FSS_VCODETPWR]        # 80 MHz VCO PWR
[C3:PSL-FSS_VCOTESTSW]        # enable/disable test input
[C3:PSL-FSS_VCOWIDESW]        # enable/disable wideband input

######################
# other FSS channels #
######################
[C3:PSL-FSS_SW1]            # frequency servo front panel switch
[C3:PSL-FSS_SW2]            # frequency servo front panel switch
[C3:PSL-FSS_INOFFSET]        # 21.5 MHz mixer input offset adjust
[C3:PSL-FSS_MGAIN]            # frequency servo common gain
[C3:PSL-FSS_FASTGAIN]        # phase correcting EOM gain
[C3:PSL-FSS_PHCON]            # 21.5 MHz phase control
[C3:PSL-FSS_RFADJ]            # 21.5 MHz oscillator output
[C3:PSL-FSS_SLOWDC]            # slow actuator voltage
[C3:PSL-FSS_MODET]            #
[C3:PSL-FSS_PHFLIP]            # 21.5 MHz 180 degree phase flip
[C3:PSL-FSS_MIXERM]            # 21.5 MHz mixer monitor
[C3:PSL-FSS_SLOWM]            # slow actuator voltage monitor
[C3:PSL-FSS_TIDALINPUT]        #
[C3:PSL-FSS_RFPDDC]            # 21.5 MHz photodetector DC output
[C3:PSL-FSS_LODET]            # detected 21.5 MHz output
[C3:PSL-FSS_PCDET]            #
[C3:PSL-FSS_FAST]            # fast actuator voltage
[C3:PSL-FSS_PCDRIVE]            # drive to the phase correcting EOM
[C3:PSL-FSS_RCTRANSPD]        # reference cavity transmission
[C3:PSL-FSS_RMTEMP]            # room temperature
[C3:PSL-FSS_RCTEMP]            # reference cavity temperature
[C3:PSL-FSS_HEATER]            # reference cavity heater power
[C3:PSL-FSS_TIDALOUT]        #
[C3:PSL-FSS_RCTLL]            # reference cavity transmitted light level
[C3:PSL-FSS_RAMP]            # slow actuator ramp, used in lock acquisition

################
# PMC channels #
################
[C3:PSL-PMC_SW1]            # PMC servo front panel switch
[C3:PSL-PMC_SW2]            # PMC servo front panel switch
[C3:PSL-PMC_MODET]
[C3:PSL-PMC_PHFLIP]            # 35.5 MHz 180 degree phase flip
[C3:PSL-PMC_PHCON]            # 35.5 MHz phase control
[C3:PSL-PMC_RFADJ]            # 35.5 MHz oscillator output
[C3:PSL-PMC_PMCERR]            # PMC error point
[C3:PSL-PMC_RFPDDC]            # 35.5 MHz photodetector DC output
[C3:PSL-PMC_LODET]            # detected 35.5 MHz output
[C3:PSL-PMC_PMCTRANSPD]        # PMC transmission
[C3:PSL-PMC_PCDRIVE]            #
[C3:PSL-PMC_PZT]            # PMC PZT voltage
[C3:PSL-PMC_INOFFSET]            # 35.5 MHz mixer input offset adjust
[C3:PSL-PMC_GAIN]            # PMC loop gain
[C3:PSL-PMC_RAMP]            # PMC PZT ramp, used in lock acquisition
[C3:PSL-PMC_BLANK]            # blanking input to the PMC PZT
[C3:PSL-PMC_PMCTLL]            # PMC transmitted light level

################
# ISS channels #
################
[C3:PSL-ISS_SW1]            # intensity servo front panel switch
[C3:PSL-ISS_SW2]            # intensity servo front panel switch
[C3:PSL-ISS_AOMRF]            # rf drive for intensity stabilization
[C3:PSL-ISS_ISERR]            # intensity servo error point
[C3:PSL-ISS_GAIN]            # intensity servo gain
[C3:PSL-ISS_ISET]            # intensity servo set point

#############################
# 16bit D/A channels - ACAV #
# 4116-card               #
#############################
[C3:PSL-ACAV_HEATER]            # analyzer cavity heater power
[C3:PSL-ACAV_SLOWDC]            # feedback to tidal input of other cavity

#############################
# 16bit A/D channels - ACAV #
# 3123-card                    #
#############################
[C3:PSL-ACAV_RCTEMP]            # reference cavity temperature
[C3:PSL-ACAV_RMTEMP]            # room temperature
[C3:PSL-ACAV_RCTRANSPD]         # analyzer cavity transmission
[C3:PSL-ACAV_RFPDDC]            # RF photodetector DC output
[C3:PSL-ACAV_PDHOUT]            # PDH servo output signal

#############################
# software channels - ACAV  #
#############################
[C3:PSL-ACAV_KP]                # pid loop p-gain
[C3:PSL-ACAV_KI]                # pid loop i-gain
[C3:PSL-ACAV_KD]                # pid loop d-gain
[C3:PSL-ACAV_LOCKEDLEVEL]       # threshold level below which pid does nothing
[C3:PSL-ACAV_TIMEOUT]           # pid loop sample time
[C3:PSL-ACAV_VERSION]           # pid loop software version
[C3:PSL-ACAV_DEBUG]             # pid loop debug messages on/off
[C3:PSL-ACAV_ENABLE]            # pid loop on/off
[C3:PSL-ACAV_SETPT]             # temperature setpoint
[C3:PSL-ACAV_SCALE]             # scaling factor

##############################
# software channels - REFCAV #
##############################
[C3:PSL-RCAV_KP]                # pid loop p-gain
[C3:PSL-RCAV_KI]                # pid loop i-gain
[C3:PSL-RCAV_KD]                # pid loop d-gain
[C3:PSL-RCAV_LOCKEDLEVEL]       # threshold level below which pid does nothing
[C3:PSL-RCAV_TIMEOUT]           # pid loop sample time
[C3:PSL-RCAV_VERSION]           # pid loop software version
[C3:PSL-RCAV_DEBUG]             # pid loop debug messages on/off
[C3:PSL-RCAV_ENABLE]            # pid loop on/off
[C3:PSL-RCAV_SETPT]             # temperature setpoint
[C3:PSL-RCAV_SCALE]             # scaling factor

##############################
# software channels - TIDAL  #
##############################
[C3:PSL-TIDAL_KP]               # pid loop p-gain
[C3:PSL-TIDAL_KI]               # pid loop i-gain
[C3:PSL-TIDAL_KD]               # pid loop d-gain
[C3:PSL-TIDAL_LOCKEDLEVEL]      # threshold level below which pid does nothing
[C3:PSL-TIDAL_TIMEOUT]          # pid loop sample time
[C3:PSL-TIDAL_VERSION]          # pid loop software version
[C3:PSL-TIDAL_DEBUG]            # pid loop debug messages on/off
[C3:PSL-TIDAL_ENABLE]           # pid loop on/off
[C3:PSL-TIDAL_SETPT]            # temperature setpoint
[C3:PSL-TIDAL_SCALE]            # scaling factor

  39   Wed Jan 27 11:25:33 2010 FrankLaserRefCavfrequency findings

Quote:

I guess if you have a frequency counter with a GPIB interface or a simple flip-flop XOR based phase/frequency discriminator, you can feed the output to a 3113 and use at as an input to a slow EPICS PID to bring the beat frequency to within range.

Actually, we need a frequency discriminator for the Green Locking so it might be good to brainstorm about this with Aidan and Koji.

 good idea -  we now know where we have to be +- acouple of 100mK as the impact of changes in room temp is not equal for both systems so we have to slightly adjust the temperature of one cavity.

Right now we have a much larger problem with the AOM. if we lock both cavities and the AOM frequency changes slowly over time (both cavities drift a bit) we get a huge amplitude modulation for the refcav, meaning a drift in power from almost 0 to 2.5V (the full range) topped by a sine-like modulation, looks like an etalon with about 10% modulation depth, every 1.4Mhz. It's not the power of the VCO, that changes only by about .2%, but you see this modulation there as well, but almost covered by the noise of the DAQ already. It's a power modulation of the diffracted beam as you can see it in the reflected and transmitted light with the same sign, so it's not pointing or so. I think it could be an impedance matching problem causing some standing wave resonator. We try to investigate it today. I will post a graph later today...

  38   Tue Jan 26 03:38:10 2010 ranaLaserRefCavfrequency findings

I guess if you have a frequency counter with a GPIB interface or a simple flip-flop XOR based phase/frequency discriminator, you can feed the output to a 3113 and use at as an input to a slow EPICS PID to bring the beat frequency to within range.

Actually, we need a frequency discriminator for the Green Locking so it might be good to brainstorm about this with Aidan and Koji.

  37   Sun Jan 24 18:35:47 2010 FrankLaserRefCavnew slow actuator values

- more insulation to the refcav chamber added

actual slow actuator values for both cavities:
refcav : 0.4385
acav : 0.2539

-> ~215MHz difference

with ~6mK/Mhz  -> refcav temp reduced by 1.3K

01/25/10  8am:

refcav temp now 67.7, slow actuator 0.304 -> reduced temp further to 67.5

01/26/10  1pm: 

refcav : 0.243
acav : 0.251

01/27/10  11am

refcav: 0.2456    temp:67.4
acav: 0.2582      temp: 40.0

01/27/10  9 pm

refcav: 0.2423    temp:67.55
acav: 0.2578      temp: 40.0

01/28/10  1am

refcav: 0.2525   temp:67.65
acav: 0.2578      temp: 40.0

01/29/10  4 pm

refcav: 0.2485   temp:67.65
acav: 0.2335      temp: 40.0

  36   Fri Jan 15 20:19:54 2010 FrankLaserRefCavboth cavities locked

after some trouble with fluctuating temperatures and a brocken cable to the vco i could finally lock both cavities. So now we can take first data.

temperatures are still fluctuating and the software loops have to be optimized for the new thermal insulation. Currently only a p-servo is running and the channel names have to be finalized...

channel names:

slow actuator signal: C:PSL-FSS_SLOWDC
room temp: C3:PEM-SENS1_TEMP

refcav:

actuator signal: C:PSL-FSS_HEATER
temp signal: C:PSL-FSS_MINCOMEAS
p: C:PSL-GEN_KP
i: C:PSL-GEN_KI
d: C:PSL-GEN_KD
transPD: C:PSL-FSS_TRANSPD

acav:

actuator signal: C3:PSL-GEN_D2A1
temp signal: C3:PSL-GEN_DAQ1
p: C3:PSL-GEN_KP
i: C3:PSL-GEN_KI
d: C3:PSL-GEN_KD
transPD: C:PSL-PMC_PMCTRANSPD

  35   Fri Jan 15 12:53:31 2010 FrankLaserRefCavlatest slow actuator values

refcav: 0.9512 @ 70.02 degC

acav: 0.9252 @ 39.996 degC

with 0.858V/GHz  ->  ~30MHz difference

1pm: changed refcav temp to 69.9 degC -> SLOWDC=0.9488

4pm: changed refcav temp to 69.5 degC -> SLOWDC=0.928 (but not final)

7pm: acav now at 0.923, refcav at 0.928  -> difference now ~6MHz

--> good temp values are ~69.5degC for the refcav and ~40degC for the acav

 

heater value refcav: ~2.983
heater value acav: ~6.245
both @21.6 degC room temp

  34   Fri Jan 15 12:28:54 2010 FrankLaserRefCavfoam cutting machine

in order to build a custom fit insulation for the cavities i've built some hot wire foam cutting machines to cut the 2" thick polyurethane thermal insulation. The wire i use is a .010 piano wire typically used for suspensions (thats what we had i the lab). Nominal current is about 1.5A at a couple of volts, so any simple power supply does the job.

Here are some pictures of the large foam cutter to cut the 48" x 48" boards into smaller peaces and the circle cutting device:

100_0476.JPG

100_0483.JPG

 

original panel size
100_0509.JPG

 

pre-cut parts:
100_0510.JPG

 

cutting into smaller segments:
100_0513.JPG

100_0516.JPG

 

surface comparison before and after hot wire cutting:
100_0504.JPG

100_0505.JPG

 

circle cutter:
100_0530.JPG

100_0535.JPG

 

some cut parts:
100_0526.JPG

 

ready to cut the rest...
100_0538.JPG

  33   Fri Jan 15 11:35:17 2010 FrankLaserRefCavrefcav update

the last couple of days we fixed a couple of thinks:

We found a wrong calibration of the temp readout of the first refcav, which gave us a wrong residual noise level compared to the other one. The absolute value was almost correct so we didn't realize this before. Now we are limited by the DAQ noise so the next step will be improving the gain settings in the temp sensor readout-box.

A big problem is the changing room temperature as soon someone opens the door even for only a couple of seconds. The temp control of the lab goes crazy and changes a couple of degrees and is oscillating half a day after that. The delay to the cavities is about 30min and a couple of deg changes of the room temp also change the cavity temperature, only a couple of tens of mK but enough to shift both cavities away from each other as the room temp couples different into both due to the different thermal insulation (insulation and time constants are different). We opened the room temp sensor and adjusted the temperature to i higher level, removed surrounding parts to get a better air flow to the sensor. We don't understand why opening the door has such a large effect on the room temp as 10 seconds of open door don't change the room temp by 2 degC or so. So if anybody has an idea plz let us know even if it seems to be stupid.

In order to improve the temp stability of the analyzer cavity we added a second layer of insulation and wrapped the whole thing in aluminum foil, see picture below. The insulation is 2" thick, except a small part at the large flanges at the end where its only 1" thick. We also changed the gain of the temp sensor readout in order to reduce the influence of DAQ noise. If nobody is working in the lab we have a stability of about 6mKpp within hours. As soon someone is working in one of the labs this changes to tens of mK for working in other labs to about a hundred mK if working in the PSL lab. We are currently working on improved servo settings...

100_0542.JPG

  32   Tue Dec 29 19:21:11 2009 FrankLaserRefCavtemperature stability

here a plot of the temperatute stability in peter's lab (upper graph) if nobody is working in the lab (christmas) and if someone opens the door to the lab (the large spikes in room temp). lower graph shows the temp of the refcav (in-loop, no out-of-loop sensor so far. Except the large spikes stability is about 4mK/pp. I will add a single ool-sensor today (will be C:PSL-FSS_RCTEMP). We used this channel so far for the second cavity but have a second temp box since christmas...

refcav_temp1.jpg

  31   Mon Dec 28 17:01:43 2009 FrankLaserRefCavthermal insulation end caps added

we added two endcaps for the first refcav and additional insulation for the ion pump and valve.

100_0456.JPG

  30   Tue Dec 22 16:15:03 2009 FrankLaserRefCavFSS settings

settings for the FSS:

common gain : 1dB

fast gain : 15dB

phase : 1.5V

VCO power : 4.5V

RF amplifier gain : 7V

  29   Tue Dec 22 00:47:40 2009 FrankLaserRefCavcavity temp stabilization

we repaired the sensors on the first refcav and  renamed and added channels to the DAQ. The channels are now:

C:PSL-FSS_MINCOMEAS  ->  temperature of first refcav (average of 4 AD590 sensors), cal. in degC

C:PSL-FSS_HEATER  ->  heater power supply control voltage (heater voltage is x10)

C3:PSL-GEN_DAQ1  ->  temperature of second refcav (one AD590), cal in degC

C3:PSL-GEN_D2A1  ->  heater power supply control voltage (heater voltage is ~ x2.5)

  28   Fri Dec 18 13:13:12 2009 FrankComputingDAQchannels for new VME-based DAQ system now available in fb1

the new channels are available as C3:PSL-GENxxxx on fb1. We have three cards installed so far:

12bit A/D channels : C3:PSL-GEN_12DAQ1 to 64

16bit A/D channels : C3:PSL-GEN_DAQ1 to 16

16bit D/A channels : C3:PSL-GEN_D2A1 to 8

the temperature of the anaylzer cavity is C3:PSL-GEN_DAQ1 (calibrated in degC)

the driving signal for the power supply for the heater is C3:PSL-GEN_D2A1 (calibrated in volts)

  27   Thu Dec 17 18:13:01 2009 FrankElectronicsRefCavpower supply remote programming

pictures taken from the existing power supply.

.100_0437.JPG

000_0005.jpg

  26   Thu Dec 17 16:57:10 2009 FrankComputingDAQnew analyzer cavity DAQ system

we set up a second, independent DAQ system for the analyzer cavity. It has one 16bit D/A card, 16bit A/D card and 12bit A/D card. The system is called "acav1" and has the ip-address 10.0.0.3, which we can access from the ATF as well via fb1 (fb1 is in peters network too). Here is a boot screen dump of the new system with a list of the new channels at the end. The new channels are now in C3, the new numbering scheme for the PSL subsystem, not in C anymore! (but all the old channels are still C as i would take to long to change all the existing medm screens...

 

                            VxWorks System Boot


Copyright 1984-1996  Wind River Systems, Inc.





CPU: Heurikon Baja4700
Version: 5.3.1
BSP version: 1.1/1
Creation date: Dec 11 1998, 10:29:37




Press any key to stop auto-boot...
 0
auto-booting...


boot device          : ei
processor number     : 0
host name            : bdl1
file name            : /usr1/epics/baja/vxWorks
inet on ethernet (e) : 10.0.0.3:ffffff00
host inet (h)        : 10.0.0.1
user (u)             : root
flags (f)            : 0x0
target name (tn)     : acav1
startup script (s)   : /usr1/epics/acav/startup.cmd

Mapping RAM base to A32 space at 0x40000000... done.
No longer waiting for sysFail to clear. D.Barker 14th Sept 1998
Attaching network interface ei0... done.
Attaching network interface lo0... done.
Loading... 795028
Starting at 0x80010000...

Mapping RAM base to A32 space at 0x40000000... done.
No longer waiting for sysFail to clear. D.Barker 14th Sept 1998
Attaching network interface ei0... done.
Attaching network interface lo0... done.
Mounting NFS file systems from host bdl1 for target acav1:
...done
Loading symbol table from bdl1:/usr1/epics/baja/vxWorks.sym ...done


 ]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]
 ]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]
 ]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]
      ]]]]]]]]]]]  ]]]]     ]]]]]]]]]]       ]]              ]]]]         (R)
 ]     ]]]]]]]]]  ]]]]]]     ]]]]]]]]       ]]               ]]]]
 ]]     ]]]]]]]  ]]]]]]]]     ]]]]]] ]     ]]                ]]]]
 ]]]     ]]]]] ]    ]]]  ]     ]]]] ]]]   ]]]]]]]]]  ]]]] ]] ]]]]  ]]   ]]]]]
 ]]]]     ]]]  ]]    ]  ]]]     ]] ]]]]] ]]]]]]   ]] ]]]]]]] ]]]] ]]   ]]]]
 ]]]]]     ]  ]]]]     ]]]]]      ]]]]]]]] ]]]]   ]] ]]]]    ]]]]]]]    ]]]]
 ]]]]]]      ]]]]]     ]]]]]]    ]  ]]]]]  ]]]]   ]] ]]]]    ]]]]]]]]    ]]]]
 ]]]]]]]    ]]]]]  ]    ]]]]]]  ]    ]]]   ]]]]   ]] ]]]]    ]]]] ]]]]    ]]]]
 ]]]]]]]]  ]]]]]  ]]]    ]]]]]]]      ]     ]]]]]]]  ]]]]    ]]]]  ]]]] ]]]]]
 ]]]]]]]]]]]]]]]]]]]]]]]]]]]]]]
 ]]]]]]]]]]]]]]]]]]]]]]]]]]]]]       Development System
 ]]]]]]]]]]]]]]]]]]]]]]]]]]]]
 ]]]]]]]]]]]]]]]]]]]]]]]]]]]       VxWorks version 5.3.1
 ]]]]]]]]]]]]]]]]]]]]]]]]]]       KERNEL: WIND version 2.5
 ]]]]]]]]]]]]]]]]]]]]]]]]]       Copyright Wind River Systems, Inc., 1984-1997

                               CPU: Heurikon Baja4700.  Processor #0.
                              Memory Size: 0x1000000.  BSP version 1.1/1.
                             WDB: Ready.

Executing startup script /usr1/epics/acav/startup.cmd ...
shellPromptSet "acav> "
value = -2146677404 = 0x800c4d64 = shellHistSize + 0x4
cd "/usr1/epics/baja"
value = 0 = 0x0
ld < iocCore
value = -2130992736 = 0x80fba1a0 = prsrv_cast_client + 0x770
ld < drvSup
value = -2131103904 = 0x80f9ef60
ld < devSup
value = -2132426208 = 0x80e5c220 = pwf_xy566 + 0x70
ld < recSup
value = -2132416016 = 0x80e5e9f0
ld < seq
value = -2130995152 = 0x80fb9830
dbLoad "default.dctsdr"
value = 0 = 0x0
cd "/usr1/epics/acav/db"
value = 0 = 0x0
dbLoadRecords "n3113.db"
value = 0 = 0x0
dbLoadRecords "n3123a.db"
value = 0 = 0x0
dbLoadRecords "n4116.db"
value = 0 = 0x0
cd "usr1/epics/baja"
value = 0 = 0x0
TSconfigure(0)
value = 1 = 0x1
iocInit "resource.def"
############################################################################
###  @(#)EPICS IOC CORE
###  @(#)Version R3.12.2patch1 Date: 1996/04/02
############################################################################
bdl1:sh: /usr1/epics/acav/db/usr1/epics/baja/resource.def: cannot open
task: 0X80fbf430 tShell
iocLogClient: unable to connect to 164.54.8.167 port 7004 because "errno = 0x33"

Unable to start log server connection watch dog
No such Resource file - resource.def
vmi4116_addr 1f045c00
entry 0 address 0x0
entry 1 address 0x0
entry 2 address 0x0
entry 3 address 0x6000
entry 4 address 0x7000
entry 5 address 0xe000
entry 6 address 0xc014
entry 7 address 0x0
entry 8 address 0x0
entry 9 address 0xff00
entry 10 address 0x0
entry 11 address 0x0
entry 12 address 0x0
entry 13 address 0x0
entry 14 address 0x2000
entry 15 address 0x2c00
Failed to set time from Unix server
0x80fbf430 (tShell): iocInit: Database Failed during Initialization
0x80fbf430 (tShell): iocInit: All initialization complete
value = 0 = 0x0
coreRelease()
############################################################################
###  @(#)EPICS IOC CORE
###  @(#)Version R3.12.2patch1 Date: 1996/04/02
############################################################################
value = 226 = 0xe2
0x80f435f0 (CA UDP): CAS: couldnt start up online notify task
OK.

Done executing startup script /usr1/epics/acav/startup.cmd
acav> dbl
C3:PSL-GEN_12DAQ1
C3:PSL-GEN_12DAQ10
C3:PSL-GEN_12DAQ11
C3:PSL-GEN_12DAQ12
C3:PSL-GEN_12DAQ13
C3:PSL-GEN_12DAQ14
C3:PSL-GEN_12DAQ15
C3:PSL-GEN_12DAQ16
C3:PSL-GEN_12DAQ17
C3:PSL-GEN_12DAQ18
C3:PSL-GEN_12DAQ19
C3:PSL-GEN_12DAQ2
C3:PSL-GEN_12DAQ20
C3:PSL-GEN_12DAQ21
C3:PSL-GEN_12DAQ22
C3:PSL-GEN_12DAQ23
C3:PSL-GEN_12DAQ24
C3:PSL-GEN_12DAQ25
C3:PSL-GEN_12DAQ26
C3:PSL-GEN_12DAQ27
C3:PSL-GEN_12DAQ28
C3:PSL-GEN_12DAQ29
C3:PSL-GEN_12DAQ3
C3:PSL-GEN_12DAQ30
C3:PSL-GEN_12DAQ31
C3:PSL-GEN_12DAQ32
C3:PSL-GEN_12DAQ33
C3:PSL-GEN_12DAQ34
C3:PSL-GEN_12DAQ35
C3:PSL-GEN_12DAQ36
C3:PSL-GEN_12DAQ37
C3:PSL-GEN_12DAQ38
C3:PSL-GEN_12DAQ39
C3:PSL-GEN_12DAQ4
C3:PSL-GEN_12DAQ40
C3:PSL-GEN_12DAQ41
C3:PSL-GEN_12DAQ42
C3:PSL-GEN_12DAQ43
C3:PSL-GEN_12DAQ44
C3:PSL-GEN_12DAQ45
C3:PSL-GEN_12DAQ46
C3:PSL-GEN_12DAQ47
C3:PSL-GEN_12DAQ48
C3:PSL-GEN_12DAQ49
C3:PSL-GEN_12DAQ5
C3:PSL-GEN_12DAQ50
C3:PSL-GEN_12DAQ51
C3:PSL-GEN_12DAQ52
C3:PSL-GEN_12DAQ53
C3:PSL-GEN_12DAQ54
C3:PSL-GEN_12DAQ55
C3:PSL-GEN_12DAQ56
C3:PSL-GEN_12DAQ57
C3:PSL-GEN_12DAQ58
C3:PSL-GEN_12DAQ59
C3:PSL-GEN_12DAQ6
C3:PSL-GEN_12DAQ60
C3:PSL-GEN_12DAQ61
C3:PSL-GEN_12DAQ62
C3:PSL-GEN_12DAQ63
C3:PSL-GEN_12DAQ64
C3:PSL-GEN_12DAQ7
C3:PSL-GEN_12DAQ8
C3:PSL-GEN_12DAQ9
C3:PSL-GEN_DAQ1
C3:PSL-GEN_DAQ10
C3:PSL-GEN_DAQ11
C3:PSL-GEN_DAQ12
C3:PSL-GEN_DAQ13
C3:PSL-GEN_DAQ14
C3:PSL-GEN_DAQ15
C3:PSL-GEN_DAQ16
C3:PSL-GEN_DAQ2
C3:PSL-GEN_DAQ3
C3:PSL-GEN_DAQ4
C3:PSL-GEN_DAQ5
C3:PSL-GEN_DAQ6
C3:PSL-GEN_DAQ7
C3:PSL-GEN_DAQ8
C3:PSL-GEN_DAQ9
C3:PSL-GEN_D2A1
C3:PSL-GEN_D2A2
C3:PSL-GEN_D2A3
C3:PSL-GEN_D2A4
C3:PSL-GEN_D2A5
C3:PSL-GEN_D2A6
C3:PSL-GEN_D2A7
C3:PSL-GEN_D2A8
value = 0 = 0x0
acav>

  25   Thu Dec 17 13:55:07 2009 FrankComputingDAQBaja4700 jumper settings and setup

in order to get the baja4700 cpu  work the jumpers have to be like this

100_0428.JPG

the description of the jumpers can be found here:

100_0431.JPG

Attachment 3: 100_0432.JPG
100_0432.JPG
  24   Fri Dec 11 19:45:15 2009 FrankLaserRefCavLHO RefCav available

there is a full LIGO RefCav available here at LHO. They don"t use it anymore and its just sitting here on the table with the pump running. Rick offered me to ship the whole thing to Caltech in order to replace the second cavity in the PSL lab. The advantage would be that we have two identical setups, a good heater and a fitting insulation. In addition we could get analog control.electronics for temperature stabilization.

  23   Sun Dec 6 14:54:43 2009 FrankLaserRefCavtemp sensor added on refcav2

added a single temp sensor on refcav2 (AD590) last week  in order to monitor the temp of refcav2. We tried to change the heater power on both cavities to match the frequencies. So far we have no tempctrl. Changes of the room temperature have too large (and different) influence on the individual cavities. So it"s almost impossible to predict the room temp changes and change the heater power to match both cavities. We need a real temp stabilization of the chambers...

the name of this sensor is C:PSL-FSS_RCTEMP

room temp next to both chambers is measured with C:PSL-FSS_RMTEMP

 

  22   Sun Dec 6 13:20:52 2009 FrankElectronicsVCOVCO tuning

measured frequency tuning vs wideband input of VCO for calibration of measured spectra. graph coming soon...

Rana: measured spectra? Has there actually been a beat frequency measured after all these years???

 no, i mean i re-measured the slope of the frequency modulation input of the VCO with a lot more points. The  coefficient (MHz/V) changes a lot over the input range from -5V to 5V (internal gain of 2). We need this to calibrate the spectrum of the feedback-signal (into the VCO) for the 2nd cavity.

vco-tuning.jpg

Attachment 2: vco-tuning.ps
vco-tuning.ps
  21   Mon Nov 30 15:01:40 2009 FrankElectronicsVCOVCO tuning

measured frequency tuning vs wideband input of VCO for calibration of measured spectra. graph coming soon...

Rana: measured spectra? Has there actually been a beat frequency measured after all these years???

  20   Tue Nov 24 11:07:30 2009 FrankLaserRefCavslow actuator settings

slow actuator settings for being resonant:

update @1pm:

refcav2: 0.4649
refcav1: 0.4469

heater refcav1: programming voltage 1.95V
heater refcav2: power supply voltage 16V

  19   Mon Nov 23 17:05:51 2009 FrankLaserRefCavheater and insulation added

added the four Minco heaters and a first layer of 1" thick foam insulation to the second refcav chamber. Pressure is down to 4e-8 torr. started heating with ~20W to bring it above room temperature.

100_0376.jpg

100_0379.jpg

100_0383.jpg

100_0391.jpg

  18   Sat Nov 21 17:27:32 2009 FrankMiscRefCavold Minco heaters

found four old Minco heaters (model HR5494-106) (from 1995) .  This type with 106 Ohms is not in their system anymore.
But corresponding to their data the maximum current for this type of heater is about 7.5A. So driving this heater with 24V would give us 5.4W of heating power beeing well below the limit. Using the standard power supply for heating refcavs we can get even more power. Due to the age (14years!) the adhesive back is not sticky anymore so i will use aluminum tape for first tests.

Model
HR5494 
Silicone Rubber heater :
4.00" x 8.00"
(101.60 mm x 203.20 mm)
Heater Shape
Effective Area : 30.3 in2
(195.484 cm2)
 
Lead Gauge : AWG 24
 
Max. current : 7.5 Amps
 
Configure this model in Kapton
Application Information Actual Current : 0.21 Amps
Operating Temperature : °C   Heat Output at Volts
Resistance in Ohms
378 1.52 Watts 0.05 W/in2 (0.0078 W/cm2)
189 3.05 Watts 0.1 W/in2 (0.0155 W/cm2)
114 5.05 Watts 0.17 W/in2 (0.0264 W/cm2)
55.1 10.45 Watts 0.34 W/in2 (0.0527 W/cm2)
37.6 15.32 Watts 0.51 W/in2 (0.0791 W/cm2)
26.3 21.9 Watts 0.72 W/in2 (0.1116 W/cm2)
13.1 43.97 Watts 1.45 W/in2 (0.2248 W/cm2)
Leadwire Length
Units Inches    Centimeters
Up to 12" included in price. Over 12", up to 80" add $3.00
Heater Backing - Based on Operating Temperature (see above)
Type Temperature Maximum Allowable
Watt Density
A No Backing -45 to 235° C    
    Installation options for unbacked heaters
    #20 Stretch Tape -45 to 200° C   1545.3 W (51 W/in2)
    #6 RTV -45 to 235° C   887.4 W (29.3 W/in2)
    Factory Vulcanized -45 to 235° C   1656.4 W (54.7 W/in2)
B #12 PSA -45 to 177° C   1292.4 W (42.7 W/in2)
D 0.003" Aluminum Foil Backing -45 to 235° C    
E Foil with Acrylic PSA -45 to 150° C   1100.4 W (36.3 W/in2)
F Foil with #12 PSA -45 to 204° C   1100.4 W (36.3 W/in2)
UL Component Recognition
Check to have heater marked for UL component recognition
Part Number (as configured):   HR5494R114L12B
  17   Thu Nov 19 18:50:33 2009 AlbertoComputingComputersElog debugging output - Down time programmed today to make changes

Quote:

We want the elog process to run in verbose mode so that we can see what's going. The idea is to track the events that trigger the elog crashes.

Following an entry on the Elog Help Forum, I added this line to the elog starting script start-elog-nodus:

./elogd -p 8080 -c /cvs/cds/caltech/elog/elog-2.7.5/elogd.cfg -D -v > elogd.log 2>&1

which replaces the old one without the part with the -v argument.

The -v argument should make the verbose output to be written into a file called elogd.log in the same directory as the elog's on Nodus.

I haven't restarted the elog yet because someone might be using it. I'm planning to do it later on today.

So be aware that:

We'll be restarting the elog today at 6.00pm PT. During this time the elog might not be accessible for a few minutes.

 I tried applying the changes but they didn't work. It seems that nodus doesn't like the command syntax.

I have to go through the problem...

The elog is up again.

 

  16   Thu Nov 19 16:06:30 2009 AlbertoComputingComputersElog debugging output - Down time programmed today to make changes

We want the elog process to run in verbose mode so that we can see what's going. The idea is to track the events that trigger the elog crashes.

Following an entry on the Elog Help Forum, I added this line to the elog starting script start-elog-nodus:

./elogd -p 8080 -c /cvs/cds/caltech/elog/elog-2.7.5/elogd.cfg -D -v > elogd.log 2>&1

which replaces the old one without the part with the -v argument.

The -v argument should make the verbose output to be written into a file called elogd.log in the same directory as the elog's on Nodus.

I haven't restarted the elog yet because someone might be using it. I'm planning to do it later on today.

So be aware that:

We'll be restarting the elog today at 6.00pm PT. During this time the elog might not be accessible for a few minutes.

  15   Wed Nov 18 20:52:39 2009 FrankLaserLaserlaser error

this morning the laser was off with the error "HT error". As the chiller was off too i think that means high temperature error, right?
So i checked everything and started the laser again. So far everything is fine and working. Any idea?

  14   Wed Nov 18 20:45:52 2009 FrankMiscRefCavleaking vaccum valve

after getting the o-ring from the drever lab for the adapter to the turbo pump to pump the second chamber i figured out that the installed CF2.75 valve is leaking.
So while pumping with an external pump everything is fine but after disconnecting the external pump the ion pump can"t handle the leak rate and the pressure increases to a level where the ion pump switches off
.
The leak is so large that you can"t disconnect the external one and quickly close the open port with a blank one until the pressure reaches the limit for the ion pump.
Because we don"t have a spare valve i decided to close this leaking port and don"t use this valve for further pumping. Instead i build an adapter to the small valve for a tiny hose (<1/4") already installed.
By using this adapter the pump rate is tiny but i still have the chance to lower the pressure below the limit of the ion pump.
By now the pressure is low enough to switch on the ion pump. The current is less than 10mA by now and decreasing, so i think the pressure should be ok tomorrow.
 

In parallel i started building the thermal insulation for that chamber.

  13   Tue Nov 17 12:01:54 2009 FrankMiscRefCavheaters for the other two chambers

IQuote:

For the heater, my main concern is on the residual thermal gradients after stabilization. The Alnis, Hansch, et al papers described how they had trouble with using multiple heaters and multiple sensors.

Our experience from the 40m is also that the single out of loop sensor (AD590) shows a much larger signal than the residual in-loop noise. This is only an issue below 10 mHz, so its not worth worrying about if it makes things take longer in the near term.

 

my hope is that if we choose the right heaters with the same heating power density that we will not have large thermal gradients. e.g. if we choose 4 heaters in total, 2 small ones and 2 large ones and we choose the right values for the resistance (you have the choice of 5 different values)  then e.g. we first could connect a small and a large one in series and then both pairs in parallel or so. I don't wanna have individual drivers for the individual heaters. That should give us a fairly even temperature distribution of the entire chamber.

So far our plan is also that we add 5 or more sensors on each layer of insulation/heating in order to measure the thermal distribution and gradient throughout the entire insulation system.

  12   Fri Nov 13 12:46:35 2009 ranaMiscRefCavheaters for the other two chambers

For the heater, my main concern is on the residual thermal gradients after stabilization. The Alnis, Hansch, et al papers described how they had trouble with using multiple heaters and multiple sensors.

Our experience from the 40m is also that the single out of loop sensor (AD590) shows a much larger signal than the residual in-loop noise. This is only an issue below 10 mHz, so its not worth worrying about if it makes things take longer in the near term.

 

  11   Fri Nov 13 11:12:38 2009 FrankLaserRefCav temp settings for both cavities

current settings for the slow actuator to bring the cavities on resonance are:

  • refcav 1 : 0.6836V
  • refcav 2 : 0.7350V

df~54MHz

  10   Fri Nov 13 10:56:32 2009 FrankThings to BuyRefCavheaters for the other two chambers

Quote:

I think its important to think of a good solution, since we may want to retrofit some other chambers.

For the electronics to drive the heater, let's make sure not to use the power supply solution that is in use now in LIGO and at the 40m. We should make sure to make the design such that the residual temperature noise from the heater is below what we expect for coating thermal noise, assuming we use Fused Silica spacer of 1m length.

This should be quite easy in principle, especially with the radiative shields on the inside.

 

 What is a "good" solution for you? A custom-made heater fitting the whole thing like the ones at the sites? They are expensive and it will take several weeks/month  to get them. The only advantage i see is that you can remove them from the chamber but do we need this feature? The other heaters are standard parts and you just stick them to the chamber, so you have a good thermal contact to it. As we only have two of the "other" chambers with a different layout for the small CF flanges i think this is good enough for our tests. For a new super-cavity chamber we can choose a different design. I think it does not make a big difference in stability. The design of the insulation is more important..

Yeah, i'm currently thinking off designing a first prototype for such a low-noise driver. As i think we will use the DAQ for temp controll thats the only part we need (and the power supply for the supply, but we have one of those). The current heaters i want to use have enough heating power when used with up to 24V, which is a good value as easy to buy. The noise can be easy as low as a couple of 10nV/sqrt(Hz) with standard parts..

  9   Thu Nov 12 20:56:38 2009 ranaThings to BuyRefCavheaters for the other two chambers

I think its important to think of a good solution, since we may want to retrofit some other chambers.

For the electronics to drive the heater, let's make sure not to use the power supply solution that is in use now in LIGO and at the 40m. We should make sure to make the design such that the residual temperature noise from the heater is below what we expect for coating thermal noise, assuming we use Fused Silica spacer of 1m length.

This should be quite easy in principle, especially with the radiative shields on the inside.

 

  8   Thu Nov 12 18:01:10 2009 FrankThings to BuyRefCavheaters for the other two chambers

in order to improve the stability of the chamber temperature the current plan is to add a heater and insulation for the second chamber. room temperature changes were about 2Kpp over the last couple of days. I already ordered flexible insulating foam for the chamber (the round parts). What we need is one or more heaters. We could somehow add half of the original heater to that chamber but i would like to go for a more final solution as we need one for the other chamber in the ATF as well. The plan is to buy standard heaters with adhesive backside and stick them to the chamber.  Price is about $70 for a 10"x10" heater (MINCO). The chamber surface is about 22"x25" in total, cut into smaller areas by the 4 vaccum tubes and 2 legs. I think we can cover most of it with a total amount of 4 to 6 heaters of different sizes.

  7   Thu Nov 12 17:45:02 2009 FrankLaserRefCavlaser crzstal temp settings for both cavities

currently the setting for the NPRO temp slider is:

  • refcav 1: 0.7341
  • refcav 2: 0.7570

changes for refcav2 are large due to the missing insulation and heater on that chamber. The LIGO refcav heaters don"t fit as the position of the connections to the main chamber tube are different (and in quantity too). changes over the last couple of hours are .7350 to .7570, which is abou 20MHz and so more than the tuning range of the VCO

  6   Thu Nov 12 17:39:49 2009 FrankElectronicsVCOtuning range

tuning range of the 80MHz VCO used for the frequency stabilization:

tuning.pdf

tuning.png

  5   Mon Nov 9 18:49:15 2009 FrankLaserRefCavresonance frequencies

Quote:

I don't know if its worth the trouble, but we do have a ~200 MHz AOM. Sam Waldman had us buy one of these for doing the OMC g-factor measurements.

 

 i think the problem is that we don't have the VCO for the FSS for 200MHz. so i think it"s easier to heat one of the cavities. the temperature required to match them should be only a couple of Kelvin difference. and by heating one of those cavities the frequency noise due to ambient temperature fluctuations might be uncorrelated as well.

i had a problem taking the data from one of the photodiodes over the weekend. We had a loose connection for the cavity1 transmitted light PD. i only checked on a scope and assumed that the (already) connected cable to the DAQ is OK as i saw some fluctuations while scanning. so i have no data from that long scan over the weekend. we repaired this today and also made the FSS for the first cavity work. now everything is working and i started a new measurement...

sweep frequency is 50uHz which is almost 3h for the full ramp from min to max

C:PSL-FSS_RMTEMP is used for the heating voltage set point (not the heater voltage!). calibration is [-273-(voltage*100)] (is still calibrated for temperature sensors)

  • cavity1 is on resonance if C:PSL-FSS_RCTRANSPD is ~1.8V
  • cavity2 is on resonance if C:PSL-PMC_TRANSPD is ~780mV

we will add the temp sensors on the cavity tank tomorrow...

  4   Mon Nov 9 02:19:45 2009 ranaLaserRefCavresonance frequencies

I don't know if its worth the trouble, but we do have a ~200 MHz AOM. Sam Waldman had us buy one of these for doing the OMC g-factor measurements.

 

  3   Sun Nov 8 14:38:56 2009 FrankLaserDAQchannels for the temperature scan

the channels used for the refcav temp scan are:

  • C:PSL-PMC_PMCTRANSPD  > transmitted light of refcav2
  • C:PSL-FSS_RCTRANSPD > transmitted light of refcav1
  • C:PSL-PMC_RMTEMP > analog programming voltage of the heater power supply

locked the laser to cavity2 using the PZT and scanning the temp of cavity1

  2   Sun Nov 8 14:35:25 2009 FrankLaserRefCavresonance frequencies

i checked the resonant frequency of both cavities in order to see if we can lock both using the existing frequency actuator (AOM) for the first one. Used the slow frequency actuator of the NPRO (temp) to scan the frequency.

refcav1 is resonant @ 0.7578V

refcav2 is resonant @ 0.5068V

assuming about 1GHz/V the resonances are about 250MHz different. So we have to use the thermal actuator on one of the cavities in order to tune it. I started a calibration scan for the heater on friday in order to set the correct heating power as the time constant is more than an hour and trial and error method would take too long...

  1   Thu Nov 5 07:07:40 2009 ranaMiscRC noiseFrank's pictures in other log

http://nodus.ligo.caltech.edu:8080/AdhikariLab/423

ELOG V3.1.3-